CPI, Inc. is acquiring JW for R470 000 in cash. CPI has 27 000 shares outstanding at a market value of R320 a share

 2. POST ACQUISITION VALUE

 CPI, Inc. is acquiring JW for R470 000 in cash. CPI has 27 000 shares outstanding at a market value of R320 a share. JW has 32 000 shares outstanding at a market price of R140 a share. Neither firm has any debt. The synergy value of the acquisition is R18 000.

 What is the value of CPI after the acquisition?


 3. NUMBER OF NEW SHARES TO BE ISSUED FOR ACQUISITION

 GM Corporation is being acquired by BKF Ltd. for R1 290 000 worth of BKF Ltd. shares. BKF has 75 000 shares outstanding at a price of R54 a share. GM has 15 000 shares outstanding with a market value of R27 a share. The synergy value of the acquisition is R37 000.

 How many new shares will be issued to complete this acquisition?


 4. AFTER MERGER EARNINGS

 Firm XY is planning on merging with Firm YZ. Firm XY will pay Firm YZ's shareholders the current value of their shares in shares of Firm XY. Firm XY currently has 39 000 shares outstanding at a market price of R40 a share. Firm YZ has 22 000 shares outstanding at a price of R17 a share. The after-merger earnings will be R78 000.

 What will the earnings per share be after the merger?
Answer #1

*AS PER HOMEWORKLIB POLICY I'VE ANSWERED 1ST QUESTION. KINDLY POST QUESTIONS SEPERATELY(ONE QUESTION AT A TIME)

Value of CPI after acquisition:

= [Market value per share of CPI × No. of CPI shares outstanding ] - cash paid +[Market value per share of JW × shares of JW outstanding] + synergy value

=(27000×320) -470000+[140×32000] +18000

=$12,668,000
Answer #2

(2)In the given case,

Acquirer =CPI,Inc

Acquiree=JW

Value of CPI after acquisition

=Value of CPI before acquisition+Value of JW before acquisition +Synergy

=27000sh×R320+32000sh×R140+R18000

=R13,138,000

(3)Here Acquirer is BKF,Acquiree is GM corporation.

BKF ltd has agreed to give it's shares worth R1,290,000 to GM corporation to complete the acquisition.

Value of share of BKF ltd=R54

So, number of shares to be issued to GM

=R1,290,000÷R54

=23889.

Therefore,BKF ltd will issue 23889 shares of its own to give a compensation worth R1,290,000 to the shareholders of GM.

(4)Here both the firms are being merged

Post merger EPS =(EATXY+EATYZ)÷n+■n

Where EAT is Earnings after tax, n is number of shares in XY,■n is number of shares to YZ.

■n=Value of YZ÷Market value per share for XY(as per the terms of agreement)

=(22000sh×R17)/R40

=9350 shares

Total number of shares post merger(n+■n)=39000+9350=48350shares

Post merger EPS

=Post merger earnings÷Post merger shares

=R78000÷48350sh

=R1.6132 per share(rounded off to four decimals)